OpenStudy (anonymous):

Explain, in complete sentences, how to factor 4x2 – 4x – 48. Explain how your factors could be checked for accuracy.

OpenStudy (anonymous):

well something like n * (x+a) * (x+b) n*x^2+n*x*b+n*a*x+n*a*b just find values that fit 4x2 – 4x – 48 for example n=4

OpenStudy (pfenn1):

\[4x^2-4x-48=4(x^2-x-12)=4(x+a)(x+b)\]where (a + b) = -1 and ab = -12. So what factors of -12 when added to each other equal -1?